# algorithmic

### v1

@@ Appendix D State evolution
$M^{z,t}_{w}$ $=\frac{1}{n_{z}-1}\sum_{i\neq j}^{n_{z}}\hat{w}^{z,t}_{i\to ij}w_{z,i}^{0}\quad$ $M^{z,t}_{v}$ $=\frac{1}{d-1}\sum_{j\neq i}^{d}\hat{v}^{z,t}_{i\to ij}v_{z,j}^{0}$ (D.1)
$Q^{z,t}_{w}$ $=\frac{1}{n_{z}-1}\sum_{i\neq j}^{n_{z}}\hat{w}^{z,t}_{i\to ij}\hat{w}^{z,t}_{i\to ij}\quad$ $Q^{z,t}_{v}$ $=\frac{1}{d-1}\sum_{j\neq i}^{d}\hat{v}^{z,t}_{i\to ij}\hat{v}^{z,t}_{i\to ij}$ (D.2)
$\Sigma^{z,t}_{w}$ $=\frac{1}{n_{z}-1}\sum_{i\neq j}^{n_{z}}\hat{\sigma}^{z,t}_{w,i\to ij}\quad$ $\Sigma^{z,t}_{v}$ $=\frac{1}{d-1}\sum_{j\neq i}^{d}\hat{\sigma}^{z,t}_{v,i\to ij}$ (D.3)
@@
$\frac{1}{d}\sum_{k}^{d}f_{kl}=f+\mathcal{O}\left(\frac{1}{\sqrt{d}}\right)=\mathbb{E}(f_{kl})+\mathcal{O}\left(\frac{1}{\sqrt{d}}\right).$ (D.4)
@@
Since the order parameters are self-averaging we replace the sum over node indices by an average over the disorder, and write their update equations by plugging in the rBP equations (A.14-A.17). At the example of $M_{w}$,
$M^{z,t}_{w}=\mathbb{E}_{w_{z}^{0},K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij}}[f_{\textrm{in}}^{w}(K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij})w_{z}^{0}].$ (D.5)
@@
$\forall w_{i},v_{j}\quad\int\mathrm{d}z_{ij}\;P^{0}_{\mathrm{out}}(z_{ij}|w^{z}_{i},v^{z}_{j})\frac{\partial g_{z}(z_{ij}|w^{z}_{i},v^{z}_{j})}{\partial w^{z}_{i}/v^{z}_{j}}=0,$ (D.6)
which in the Bayes-optimal case $P^{0}_{\mathrm{out}}(z_{ij}|w^{z}_{i},v^{z}_{j})=e^{g_{z}(z_{ij}|w^{z}_{i},v^{z}_{j})}$ follows directly from normalization. For a discussion of when this is satisfied, refer to Lesieur et al. (2017), p.34. We do the mean and variance calculation first at the example of $J^{w,t}_{z,i\to ij}$. The mean is
$\mathbb{E}(J^{w,t}_{z,i\to ij})$ $=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k\neq i}^{d}\int\mathrm{d}z_{ik}\;P^{0}_{\mathrm{out}}(z_{ik}|w_{z,i}^{0}\;v_{z,k}^{0})S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}$ (D.7)
$=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k\neq i}^{d}\int\mathrm{d}z_{ik}\;P^{0}_{\mathrm{out}}(z_{ik}|0)\left[1+\lambda^{0}_{z}\frac{w_{z,i}^{0}\;v_{z,k}^{0}}{\sqrt{n_{z}}}S^{0,z}_{ik}+\mathcal{O}(\frac{1}{n_{z}})\right]S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}$ (D.8)
$=\frac{\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}w_{z,i}^{0}\;\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[\frac{1}{n_{z}}\sum_{k\neq i}^{d}\hat{v}^{z,t}_{k\to ik}v_{z,k}^{0}\right]+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (D.9)
$=\frac{\alpha_{z}\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M^{z,t}_{v}w_{z,i}^{0}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}}).$ (D.10)
@@
$\frac{1}{\hat{\Delta}^{z}}=\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[S^{0,z}_{ik}S^{z}_{ik}\right]$ (D.11)
@@
$\mathrm{Var}(J^{w,t}_{z,i\to ij})$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sum_{k,l\neq i}^{d}\mathbb{E}_{P^{0}_{\mathrm{out}}(z_{ik}|w_{z,i}^{0}\;v_{z,k}^{0})}\mathbb{E}_{P^{0}_{\mathrm{out}}(z_{il}|w_{z,i}^{0}\;v_{z,l}^{0})}\left[S^{z}_{ik}S^{z}_{il}\hat{v}^{z,t}_{k\to ik}\hat{v}^{z,t}_{l\to il}\right]-\mathbb{E}(J^{w,t}_{z,i\to ij})^{2}$ (D.12)
$=\frac{\lambda_{z}^{2}}{n_{z}}\sum_{k\neq i}^{d}\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[\left(S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}\right)^{2}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})\right]+\mathcal{O}(\frac{1}{d})$ (D.13)
$=\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[\frac{1}{n_{z}}\sum_{k\neq i}^{d}\hat{v}^{z,t}_{k\to ik}\hat{v}^{z,t}_{k\to ik}\right]+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (D.14)
$=\frac{\alpha_{z}\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{t}_{v}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (D.15)
@@
$\frac{1}{\tilde{\Delta}^{z}}=\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[S^{z}_{ik}S^{z}_{ik}\right].$ (D.16)
@@
$\mathbb{E}(K^{w,t}_{z,i\to ij})$ $=\frac{\alpha_{z}\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t}_{v}-\alpha_{z}\lambda_{z}^{2}\bar{R}^{z}(Q^{z,t}_{v}+\Sigma^{z,t}_{v})+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (D.17)
$\mathrm{Var}(K^{w,t}_{z,i\to ij})$ $=\mathcal{O}(\frac{1}{\sqrt{n_{z}}}),$ (D.18)
where we used approximate independence of $R^{z}_{ik}$ and $((\hat{v}^{z,t}_{k\to ik})^{2}+\hat{\sigma}^{z,t}_{v,k\to ik})$ as before for $S^{z}_{ij}$ and defined using self-averaging (D.4)
$\bar{R}^{z}=\mathbb{E}_{P^{0}_{\mathrm{out}}(z_{ik})}(R^{z}_{ik})=\frac{1}{d}\sum_{k\neq i}^{d}R^{z}_{ik}+\mathcal{O}(\frac{1}{\sqrt{d}}).$ (D.19)
@@
$\mathbb{E}(J^{v,t}_{z,j\to ij})$ $=\frac{\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M^{z,t-1}_{w}v_{z,j}^{0}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (D.20)
$\mathrm{Var}(J^{v,t}_{z,j\to ij})$ $=\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t-1}_{w}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (D.21)
$\mathbb{E}(K^{v,t}_{z,j\to ij})$ $=\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t-1}_{w}-\lambda_{z}^{2}\bar{R}^{z}(Q^{z,t-1}_{w}+\Sigma^{z,t-1}_{w})+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (D.22)
$\mathrm{Var}(K^{v,t}_{z,j\to ij})$ $=\mathcal{O}(\frac{1}{\sqrt{n_{z}}}).$ (D.23)
Due to the exchange of node and disorder averages the means and variances are independent of the $i,j$ indices, so that we drop them. Also it does not make a difference with the truncation at $\mathcal{O}(\frac{1}{\sqrt{n}})$ whether the $i=j$ term is included in the marginal or not. Equipped with the statistics of the source terms, plugging the rBP equations (A.14)-(A.17) into the order parameter definitions (D.1)-(D.3) and using self-averaging as in the example (D.5), we obtain the state evolution equations:
$M^{z,t}_{w}$ $=\mathbb{E}_{w_{z}^{0},J^{w,t}_{z}}\left[f^{w}_{\mathrm{in}}(K^{w,t}_{z},J^{w,t}_{z})\;w_{z}^{0}\right]$ (D.24)
$M^{z,t}_{v}$ $=\mathbb{E}_{(v_{X}^{0},v_{Y}^{0}),J^{v,t}_{X},J^{v,t}_{Y}}\left[f^{v^{z}}_{\mathrm{in}}(K^{v,t}_{X},K^{v,t}_{Y},J^{v,t}_{X},J^{v,t}_{Y})\;v_{z}^{0}\right]$ (D.25)
$Q^{z,t}_{w}$ $=\mathbb{E}_{w_{z}^{0},J^{w,t}_{z}}\left[f^{w}_{\mathrm{in}}(K^{w,t}_{z},J^{w,t}_{z})^{2}\right]$ (D.26)
$Q^{z,t}_{v}$ $=\mathbb{E}_{(v_{X}^{0},v_{Y}^{0}),J^{v,t}_{X},J^{v,t}_{Y}}\left[f^{v^{z}}_{\mathrm{in}}(K^{v,t}_{X},K^{v,t}_{Y},J^{v,t}_{X},J^{v,t}_{Y})^{2}\right]$ (D.27)
$\Sigma^{z,t}_{w}$ $=\mathbb{E}_{w_{z}^{0},J^{w,t}_{z}}\left[\frac{\partial f^{w}_{\mathrm{in}}}{\partial J}(K^{w,t}_{z},J^{w,t}_{z})\right]$ (D.28)
$\Sigma^{z,t}_{v}$ $=\mathbb{E}_{(v_{X}^{0},v_{Y}^{0}),J^{v,t}_{X},J^{v,t}_{Y}}\left[\frac{\partial f^{v^{z}}_{\mathrm{in}}}{\partial J_{1/2}}(K^{v,t}_{X},K^{v,t}_{Y},J^{v,t}_{X},J^{v,t}_{Y})^{2}\right]$ (D.29)
@@
$J^{w,t}_{z}$ $\sim\mathcal{N}\left(\frac{\alpha_{z}\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M^{z,t}_{v}w_{z}^{0},\;\frac{\alpha_{z}\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t}_{v}\right)$ (D.30)
$J^{v,t}_{z}$ $\sim\mathcal{N}\left(\frac{\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M^{z,t-1}_{w}v_{z}^{0},\;\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t-1}_{w}\right)$ (D.31)
$K^{w,t}_{z}$ $=\alpha_{z}\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t}_{v}-\alpha_{z}\lambda_{z}^{2}\bar{R}^{z}(Q^{z,t}_{v}+\Sigma^{z,t}_{v})$ (D.32)
$K^{v,t}_{z}$ $=\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t-1}_{w}-\lambda_{z}^{2}\bar{R}^{z}(Q^{z,t-1}_{w}+\Sigma^{z,t-1}_{w}).$ (D.33)
Note that the distributions of $J^{w,t}_{z},J^{v,t}_{z}$ still depend on $w_{z}^{0},v_{z}^{0}$, therefore the average is performed also over the priors in (D.26) - (D.29). In these 12 equations, all variables are scalars, giving the low-dimensional effective description of the relaxed BP as well as the AMP dynamics.

### v2

@@ Appendix D State evolution
$M^{z,t}_{w}$ $=\frac{1}{n_{z}-1}\sum_{i\neq j}^{n_{z}}\hat{w}^{z,t}_{i\to ij}w_{z,i}^{0}\quad$ $M^{z,t}_{v}$ $=\frac{1}{d-1}\sum_{j\neq i}^{d}\hat{v}^{z,t}_{i\to ij}v_{z,j}^{0}$ (71)
$Q^{z,t}_{w}$ $=\frac{1}{n_{z}-1}\sum_{i\neq j}^{n_{z}}\hat{w}^{z,t}_{i\to ij}\hat{w}^{z,t}_{i\to ij}\quad$ $Q^{z,t}_{v}$ $=\frac{1}{d-1}\sum_{j\neq i}^{d}\hat{v}^{z,t}_{i\to ij}\hat{v}^{z,t}_{i\to ij}$ (72)
$\Sigma^{z,t}_{w}$ $=\frac{1}{n_{z}-1}\sum_{i\neq j}^{n_{z}}\hat{\sigma}^{z,t}_{w,i\to ij}\quad$ $\Sigma^{z,t}_{v}$ $=\frac{1}{d-1}\sum_{j\neq i}^{d}\hat{\sigma}^{z,t}_{v,i\to ij}$ (73)
@@
$\frac{1}{d}\sum_{k}^{d}f_{kl}=f+\mathcal{O}\left(\frac{1}{\sqrt{d}}\right)=\mathbb{E}(f_{kl})+\mathcal{O}\left(\frac{1}{\sqrt{d}}\right).$ (74)
@@
Since the order parameters are self-averaging we replace the sum over node indices by an average over the disorder, and write their update equations by plugging in the rBP equations (45-48). At the example of $M_{w}$,
$M^{z,t}_{w}=\mathbb{E}_{w_{z}^{0},K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij}}[f_{\textrm{in}}^{w}(K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij})w_{z}^{0}].$ (75)
@@
$\forall w_{i},v_{j}\quad\int\mathrm{d}z_{ij}\;P^{0}_{\mathrm{out}}(z_{ij}|w^{z}_{i},v^{z}_{j})\frac{\partial g_{z}(z_{ij}|w^{z}_{i},v^{z}_{j})}{\partial w^{z}_{i}/v^{z}_{j}}=0,$ (76)
which in the Bayes-optimal case $P^{0}_{\mathrm{out}}(z_{ij}|w^{z}_{i},v^{z}_{j})=e^{g_{z}(z_{ij}|w^{z}_{i},v^{z}_{j})}$ follows directly from normalization. For a discussion of when this is satisfied, refer to [17], p.34. We do the mean and variance calculation first at the example of $J^{w,t}_{z,i\to ij}$. The mean is
$\mathbb{E}(J^{w,t}_{z,i\to ij})$ $=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k\neq i}^{d}\int\mathrm{d}z_{ik}\;P^{0}_{\mathrm{out}}(z_{ik}|w_{z,i}^{0}\;v_{z,k}^{0})S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}$ (77)
$=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k\neq i}^{d}\int\mathrm{d}z_{ik}\;P^{0}_{\mathrm{out}}(z_{ik}|0)\left[1+\lambda^{0}_{z}\frac{w_{z,i}^{0}\;v_{z,k}^{0}}{\sqrt{n_{z}}}S^{0,z}_{ik}+\mathcal{O}(\frac{1}{n_{z}})\right]S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}$ (78)
$=\frac{\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}w_{z,i}^{0}\;\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[\frac{1}{n_{z}}\sum_{k\neq i}^{d}\hat{v}^{z,t}_{k\to ik}v_{z,k}^{0}\right]+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (79)
$=\frac{\alpha_{z}\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M^{z,t}_{v}w_{z,i}^{0}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}}).$ (80)
@@
$\frac{1}{\hat{\Delta}^{z}}=\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[S^{0,z}_{ik}S^{z}_{ik}\right]$ (81)
@@
$\mathrm{Var}(J^{w,t}_{z,i\to ij})$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sum_{k,l\neq i}^{d}\mathbb{E}_{P^{0}_{\mathrm{out}}(z_{ik}|w_{z,i}^{0}\;v_{z,k}^{0})}\mathbb{E}_{P^{0}_{\mathrm{out}}(z_{il}|w_{z,i}^{0}\;v_{z,l}^{0})}\left[S^{z}_{ik}S^{z}_{il}\hat{v}^{z,t}_{k\to ik}\hat{v}^{z,t}_{l\to il}\right]-\mathbb{E}(J^{w,t}_{z,i\to ij})^{2}$ (82)
$=\frac{\lambda_{z}^{2}}{n_{z}}\sum_{k\neq i}^{d}\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[\left(S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}\right)^{2}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})\right]+\mathcal{O}(\frac{1}{d})$ (83)
$=\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[\frac{1}{n_{z}}\sum_{k\neq i}^{d}\hat{v}^{z,t}_{k\to ik}\hat{v}^{z,t}_{k\to ik}\right]+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (84)
$=\frac{\alpha_{z}\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{t}_{v}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (85)
@@
$\frac{1}{\tilde{\Delta}^{z}}=\mathbb{E}_{P^{0}_{\mathrm{out}}(z|0)}\left[S^{z}_{ik}S^{z}_{ik}\right].$ (86)
@@
$\mathbb{E}(K^{w,t}_{z,i\to ij})$ $=\frac{\alpha_{z}\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t}_{v}-\alpha_{z}\lambda_{z}^{2}\bar{R}^{z}(Q^{z,t}_{v}+\Sigma^{z,t}_{v})+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (87)
$\mathrm{Var}(K^{w,t}_{z,i\to ij})$ $=\mathcal{O}(\frac{1}{\sqrt{n_{z}}}),$ (88)
where we used approximate independence of $R^{z}_{ik}$ and $((\hat{v}^{z,t}_{k\to ik})^{2}+\hat{\sigma}^{z,t}_{v,k\to ik})$ as before for $S^{z}_{ij}$ and defined using self-averaging (74)
$\bar{R}^{z}=\mathbb{E}_{P^{0}_{\mathrm{out}}(z_{ik})}(R^{z}_{ik})=\frac{1}{d}\sum_{k\neq i}^{d}R^{z}_{ik}+\mathcal{O}(\frac{1}{\sqrt{d}}).$ (89)
@@
$\mathbb{E}(J^{v,t}_{z,j\to ij})$ $=\frac{\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M^{z,t-1}_{w}v_{z,j}^{0}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (90)
$\mathrm{Var}(J^{v,t}_{z,j\to ij})$ $=\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t-1}_{w}+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (91)
$\mathbb{E}(K^{v,t}_{z,j\to ij})$ $=\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t-1}_{w}-\lambda_{z}^{2}\bar{R}^{z}(Q^{z,t-1}_{w}+\Sigma^{z,t-1}_{w})+\mathcal{O}(\frac{1}{\sqrt{n_{z}}})$ (92)
$\mathrm{Var}(K^{v,t}_{z,j\to ij})$ $=\mathcal{O}(\frac{1}{\sqrt{n_{z}}}).$ (93)
Due to the exchange of node and disorder averages the means and variances are independent of the $i,j$ indices, so that we drop them. Also it does not make a difference with the truncation at $\mathcal{O}(\frac{1}{\sqrt{n}})$ whether the $i=j$ term is included in the marginal or not. Equipped with the statistics of the source terms, plugging the rBP equations (45)-(48) into the order parameter definitions (71)-(73) and using self-averaging as in the example (75), we obtain the state evolution equations whose overall form again always follows a generic structure [14, 43]:
$M^{z,t}_{w}$ $=\mathbb{E}_{w_{z}^{0},J^{w,t}_{z}}\left[f^{w}_{\mathrm{in}}(K^{w,t}_{z},J^{w,t}_{z})\;w_{z}^{0}\right]$ (94)
$M^{z,t}_{v}$ $=\mathbb{E}_{(v_{X}^{0},v_{Y}^{0}),J^{v,t}_{X},J^{v,t}_{Y}}\left[f^{v^{z}}_{\mathrm{in}}(K^{v,t}_{X},K^{v,t}_{Y},J^{v,t}_{X},J^{v,t}_{Y})\;v_{z}^{0}\right]$ (95)
$Q^{z,t}_{w}$ $=\mathbb{E}_{w_{z}^{0},J^{w,t}_{z}}\left[f^{w}_{\mathrm{in}}(K^{w,t}_{z},J^{w,t}_{z})^{2}\right]$ (96)
$Q^{z,t}_{v}$ $=\mathbb{E}_{(v_{X}^{0},v_{Y}^{0}),J^{v,t}_{X},J^{v,t}_{Y}}\left[f^{v^{z}}_{\mathrm{in}}(K^{v,t}_{X},K^{v,t}_{Y},J^{v,t}_{X},J^{v,t}_{Y})^{2}\right]$ (97)
$\Sigma^{z,t}_{w}$ $=\mathbb{E}_{w_{z}^{0},J^{w,t}_{z}}\left[\frac{\partial f^{w}_{\mathrm{in}}}{\partial J}(K^{w,t}_{z},J^{w,t}_{z})\right]$ (98)
$\Sigma^{z,t}_{v}$ $=\mathbb{E}_{(v_{X}^{0},v_{Y}^{0}),J^{v,t}_{X},J^{v,t}_{Y}}\left[\frac{\partial f^{v^{z}}_{\mathrm{in}}}{\partial J_{1/2}}(K^{v,t}_{X},K^{v,t}_{Y},J^{v,t}_{X},J^{v,t}_{Y})^{2}\right]$ (99)
@@
$J^{w,t}_{z}$ $\sim\mathcal{N}\left(\frac{\alpha_{z}\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M^{z,t}_{v}w_{z}^{0},\;\frac{\alpha_{z}\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t}_{v}\right)$ (100)
$J^{v,t}_{z}$ $\sim\mathcal{N}\left(\frac{\lambda_{z}\lambda^{0}_{z}}{\hat{\Delta}^{z}}M^{z,t-1}_{w}v_{z}^{0},\;\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t-1}_{w}\right)$ (101)
$K^{w,t}_{z}$ $=\alpha_{z}\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t}_{v}-\alpha_{z}\lambda_{z}^{2}\bar{R}^{z}(Q^{z,t}_{v}+\Sigma^{z,t}_{v})$ (102)
$K^{v,t}_{z}$ $=\frac{\lambda_{z}^{2}}{\tilde{\Delta}^{z}}Q^{z,t-1}_{w}-\lambda_{z}^{2}\bar{R}^{z}(Q^{z,t-1}_{w}+\Sigma^{z,t-1}_{w}).$ (103)
Note that the distributions of $J^{w,t}_{z},J^{v,t}_{z}$ still depend on $w_{z}^{0},v_{z}^{0}$, therefore the average is performed also over the priors in (96) - (99). In these 12 equations, all variables are scalars, giving the low-dimensional effective description of the relaxed BP as well as the AMP dynamics.